# Nuclear Astrophysics

### v1

@@ 1 Introduction
The field of Nuclear Astrophysics is situated between the two disciplines of nuclear physics and of astrophysics. Each of these includes experimental/observational and theoretical branches. The foundations of this field have been laid out in pioneering papers by Burbidge et al [56] and Cameron [59, 58], and later written up in textbooks such as Donald Clayton’s “Principles of stellar evolution and nucleosynthesis” [79], then “Cauldrons in the Cosmos” by Claus Rolfs and William Rodney [407], and the more-recent “Nuclear physics of stars” by Christian Iliadis [216] and “The synthesis of the elements” by Giora Shaviv [432]. Each of these take a different look at the complex interplay that occurs between the microscopic quantum processes that govern the nuclear reaction processes and the macroscopic processes of transporting energy and materials that characterize the conditions in stellar interiors. This results in considerable uncertainty for the analysis and determination of nuclear reaction processes in cosmic sites. On the other hand we understand enough about the origin of chemical elements to describe and teach this astrophysical and nuclear-physics problem [see above-cited textbooks and, for example, 56, 80, 491, 201, 23, 257, 88, 20, and other review articles that have been written over seven decades]). Often herein the descriptions of cosmic nucleosynthesis processes are inter-woven with our plausible understanding of its cosmic sources, thus mixing better consolidated with more-uncertain aspects. For astrophysicists, a description of cosmic sources which includes nuclear transformations remains a challenge, while nuclear physicists are challenged to determine nuclear reaction processes at conditions that are typically not available in the laboratory. The data base for this science are the measured outcomes of nuclear-reaction experiments, and the observations of isotopic variety and abundances in material samples or objects throughout the universe. Combining the characteristics of nuclear reactions with those of cosmic sites and their dynamics, nuclear astrophysics exploits astronomical data in attempts to rule out the invalid components of the complex and nested modeling that represents our understanding of the compositional evolution of matter in the universe.
@@
Conferences and workshops are of great importance in such a diverse field of nuclear astrophysics. They are exemplified by the series of conferences of "Nuclei in the Cosmos" (since 1990, no. XVIII held recently111see https://indico.icc.ub.edu/event/341/ for the most-recent NIC in 2025) and of "Nuclear Physics in Astrophysics" (since 2003, no. XI held recently222see https://events.hifis.net/event/540/ for the most recent NPA in 2024), typically held every 2 years. Gathering 200 people or more, these conferences may appear to novices as either overloaded with specialist discussions in sub-fields, or too shallow to address physics issues of interest with the required depth. Yet, they are essential to mediate among sub-fields, and are complemented by workshops addressing specific topics and questions in the field. International community networking projects such as IReNA (USA and international), JINA/CENAM (USA), ChETEC/ChETEC-Infra (EU/UK), UKAKUREN/JaFNA (Japan) and INAC (China) provide an umbrella for the diverse projects and for exchanges of tools; these have an important role to maintain the exchanges among diverse research groups. The experimental challenges ahead for this field are addressed by white papers that lobby for this science with the various funding agencies, see for example Lewitowicz and Widmann [292] and Schatz et al [425]. Nuclear astrophysics as a whole appears too broad a field to be addressed in a single review paper, due to this interdisciplinary nature. Partial reviews [e.g. 517] are useful to help appreciating the interplay between achievements, uncertainties, and tacit assumptions in each of the aspects of nuclear aspects.
@@ 2.1 General considerations
Only a subset of reactions, anticipated for a specific environment, are typically considered in astrophysical models. For hydrogen burning, the different branches of the pp chains [415] and the cyclic pattern of the four CNO cycles [514] are an early example of identifying reaction sequences that plausibly operate in stars. Similarly, the nuclear-reaction paths that may occur in cosmic sites are commonly organized in processes, or reaction patterns describing a reaction flow. These share a common environmental characteristic such as temperatures, densities, and abundance of reaction agents, and link a specific set of reactions. Regimes of reactions that are expected to occur under such characteristic conditions are illustrated in Figure 1 by white arrows, the processes of nucleosynthesis that are most-commonly discussed. These have been presented in the early period of nuclear astrophysics [56], and later a variety of other such processes and paths were proposed from different reasonings and scenarios [e.g. 196, 490, 527, 212, 525, 209, 423, 392, 376, 496, 378, 73, 74, (these will be discussed below in their respective context)]. Shown are the sequences of $\alpha$ capture reactions ($\alpha$ process) that may illustrate production of 56Ni and iron group elements from light nuclei, the continuation of heavier-nuclei production through slow neutron capture / $\beta$ decay sequences (s process) along the valley of stable nuclei (black squares), the alternate neutron capture path of rapid neutron captures (r process) far on the neutron-rich side, a rapid proton capture path (rp process) on the proton-rich side, and a horizontal production of proton-rich nuclei from stability through a p or $\gamma$ process, i.e. triggered by proton releases from ($\alpha$,p) or ($\gamma$,p) reactions on neutron-deficient oxygen or neon isotopes. Not shown are the generic $\beta$-decay reaction flows that drive unstable isotopes towards the valley of stability (i.e., forming diagonal lines in this Figure), driven only by the $\beta$-decay lifetimes of each isotope and thus (largely) independent from environmental characteristics. More such processes are being discussed in nuclear astrophysics as one aims to characterize a coherent nuclear-reaction network that may operate in specific cosmic sources. So, an intermittent n process [46, 378] and the intermediate i process [60, 73] are discussed to describe reactions at neutron intensities between the extremes characterizing the r and s process. Then, a $\nu$ process [527, 437] with its variants is used to characterize neutrino-driven nuclear reactions at high neutrino fluxes as they occur in the cores of gravitational collapse supernovae, and may characteristically alter the nucleosynthesis to produce specific rare isotopes that cannot be understood from the neutron capture nor $\gamma$ processes, such as 180Ta, 138La, or 19F. Also, reaction flows with electron captures and successive $\beta$ decays or vice versa through URCA pairs of nuclei define an URCA process [34]; not altering the composition, the URCA process establishes an energy leak through neutrino production, which affects burning-site conditions. These latter processes illustrate that details of nuclear reactions in cosmic sites are often more complex, and cannot all be shown in this Figure for clarity. This overview Figure 1 provides a valuable tool, to first order, for such explorations in order to find the most-relevant regions where nuclear-reaction properties matter for cosmic nucleosynthesis. It therefore is a key figure for the communication between nuclear physicists and astrophysicists.
@@
Uncertainties on reduced Coulomb repulsion in nuclear reactions under cosmic environments result from screening of the nuclear charge because of free electrons clustering around the positive charges. Particle beams in accelerator experiments are fundamentally different in their electron environments of the reaction compared to mixtures of high-temperature plasma in cosmic sites. Screening corrections are required for each of these two cases, and are different [see review by 12]. In the laboratory, potential extremes related to the screening towards the lowest energies motivate experiments studying reactions below keV energies between light nuclei such as 3He and deuterium towards 4He fusion [65, 485, 518].
@@
Experimental studies of nuclear physics for astrophysics include many indirect techniques, such as breaking up nuclei in heavy-ion collisions, or measuring surrogate reactions with similarities to the reactions and nuclear characteristics in question [see 517, for details].
@@ 2.2.1 Experimental Techniques and Methods
After the build-up, the subsequent 12C($\alpha,\gamma$)16O reaction in stellar helium burning is often called the holy grail reaction because it determines the cosmic abundance ratio between the two isotopes that are key for organic life, 12C and 16O. Numerous measurements have been performed to determine the reaction cross section towards lower energies. However, the interpretation of the data for a reliable extrapolation towards the Gamow range around 300 keV remains a challenge. The main difficulty is in the treatment of the pronounced interference effects between the low-energy tails of high-energy resonances, direct capture contributions and the high-energy tails of sub-threshold states. More specifically, the interpretation of the reaction cross-section [see 104, 103, for recent reviews] is based on contributions from a number of broad interfering $1^{-}$ and $2^{+}$ resonances in the compound nucleus $^{16}O$ as well as an E2 direct capture to the ground state of $^{16}O$. The $\gamma$ transition therefore shows an a E1 or a E2 multipolarity, with both components adding to the total cross section. Both components are predicted to have comparable strength due to the suppression of the E1 term by the breaking of isospin symmetry by Coulomb interaction as discussed by deBoer et al [103] To determine the cross-section data and map the resonance contributions interfering with bound state components extensive angular distribution in 12C($\alpha,\gamma$)16O experiments have been performed over a wide energy range.
@@ 2.5 Beta decay and neutrino reactions
In this context, the process of neutrino oscillations may alter the neutrino flavors between the neutrino source and the region where interactions with nuclei occur. This has been recognized and studied in recent years [e.g. 255] [see recent review by 145].
@@ 2.6 Nuclear theory complements
The original description of nuclear binding energies (through determining their masses) was made as an ensemble (macroscopic) effect through the liquid-drop model [511]. Since then, there had been an impressive evolution of theories [see 342, for a personal review]. A bottom-up approach starting from the nucleon-nucleon forces (microscopic approach) was pursued, in parallel to efforts aiming at an establishment of an empiric energy density functional (macroscopic approach) [500, 412]. Nuclear mass models are a compromise between empirical descriptions of macroscopic (and more-easily measurable) components and microscopic descriptions of nucleon-nucleon interactions deriving from fundamental theory or from empirical prescriptions. Nuclear theory attempts to advance models in specific regions of interest, while astrophysical nucleosynthesis models require consistent treatment of neutron captures and $\beta$ decays across the entire path of reactions; therefore, the FRDM baseline from 1995 is still widely used, as $\beta$ decays have been evaluated for a large number of nuclei here. Nuclear mass models evolved from the macroscopic finite range liquid drop model (FRDM) [345] to Hartree-Fock Bogoliubov (HFB) model versions of different sophistication levels, adding corrections to nucleon-nucleon forces with shell and deformation effects to a Fermi gas model for nuclei, among others [see 183, for testing model alternatives]. Towards the higher range of nuclear masses, nuclear deformations play a large role. The additional degrees of freedom for nuclear rotation and oscillations provide a larger phase space, thus making predictions of effective nuclear binding quite uncertain. Experimental results are analyzed and fitted to descriptions of mass models, thus determining model parameters; often models have a large number ($\simeq$ 30) of parameters determined in this way. As a measure, the 2-neutron separation energies $S_{2n}$ between models and measurements (where available) still show an RMS difference of 800 keV (see Figure 5), while reliable astrophysical r-process calculations would require a precision of order 100 keV [210]. Weak-interaction rates can be estimated theoretically in the QRPA approach [419]. But correspondingly, also the weak-interaction processes in atomic nuclei ($\beta$-decay rates, radioactive lifetimes) cannot be reliably calculated in theoretical models, if the details of shells and nuclear structure are not covered; shell models have proven to provide more realistic rates, in better agreement with experimental values [see 280, 186, for discussions of weak interactions in cosmic environments]. As agreements with experimental values are unsatisfactory, phenomenological models are widely used. This also extends to the determination of fission barriers, correspondingly [133, 343].
@@ 3.1 Stellar interiors
Once hydrogen in the core becomes exhausted, gravitational contraction results in a temperature increase until the next nuclear fuel 4He can undergo fusion reactions to liberate nuclear binding energy again. At the same time, hydrogen burning from then on proceeds in a shell [^5] Note that only $\simeq$ 10% of the total hydrogen will have been consumed during the main-sequence evolution., continuing throughout evolution in parallel to the nuclear processes in the stellar core (see Figure 7). The radiation pressure caused by nuclear energy release in the shell causes expansion of the outer zones of the star, while the core contracts due to a lack of energy production, removing the required thermal pressure for hydrostatic stability, and thus further expanding the envelope as a reaction to the contraction of the core. This is called the giant phase in stellar evolution [see 79, for a textbook introduction to stellar evolution]. Due to the smaller energy gain per 12C produced (7.3 MeV), the total helium-burning time is not quite as long as the main-sequence time; for the case of our Sun it is $\simeq$ 10% of the main sequence phase. This seems sufficiently long for most stars to still achieve and maintain hydrostatic equilibrium, rearranging temperatures and densities to the convective and radiative energy transport processes within the star and the nuclear energy gain in the core balancing radiative and wind losses at the surface.
@@
The presupernova evolution of massive stars in their late phases is a challenging topic in astrophysics. The above-discussed processes of mixing determine how evolution proceeds and modifies the structure of the star and its hydrostatic balance [see 513, for a recent broad-impact study]. Many different scenarios can be inferred from simulations, as mixing prescriptions and other parameters are varied. Examples are: burning shells progressing further inwards (instead of the more-plausible trend to proceed along the availability of unburnt fuel at larger distance from the core) further inwards with time [235, 237], or merge [404, 221], or hot-bottom burning [249]. Even though agreement is reached on the fundamental microscopic processes, their implementation in different numerical codes may provide different results, depending on how these processes are approximated and numerically stabilized within each code. Figure 7 illustrates an example; alternative codes for stellar evolution are maintained, e.g., by the Geneva [421, 129, 173, 336, 190] and Rome [70, 71, 295] groups. Astronomical observations may provide clarifications of such issues concerning stellar interiors (see Section 4 on Nuclear Astronomy below).
@@
Stars are often formed in groups, and multiple stars may be created as single entities yet close enough to have significant mutual interactions that affects their evolution. The impact of a companion star on stellar evolution had been described long ago [215]. The rather obvious effect of a companion to enhance mass loss and thus stripping the envelope has been modeled [529, 385, 463]. This is understood as a path towards explaining core-collapse supernovae which do not show a hydrogen envelope, and are thus observationally known as supernova types Ib and Ic. More difficult are successive periods of mass exchange in different directions within a binary system, or phases of a common envelope [e.g. 504]. Thus, binary interactions and their impacts have developed into a major field of stellar evolution [215, 480, 385, 418, 99, 224, 482, 282, 319, 426]. This is mainly addressed by theory and simulations, because it is difficult to obtain and exploit observational constraints on models of binary interactions, as observational phenomena cannot easily by attributed to binary interaction physics as discriminated from instabilities in single-star evolution.
@@
Compact stars are stabilized by degeneracy pressure from electrons (the white dwarf stars), and by the complex pressure behavior of dense nuclear matter called its equation of state (the neutron stars). White dwarfs are the compact remnants of stars after the giant stage and strong mass losses therein [see 420, for a recent review]. Typical mass and size values are 0.16 M and 0.01 R, respectively. Matter is confined to a rather small volume in compact stars. Electrons move freely as a Fermi gas within the charge-neutral star which is bound by gravity that is caused mainly by the nucleons. As thermal energies become small and approach zero for the electrons, their kinetic energy level occupancy obeys quantum constrains and the Pauli exclusion principle for identical states. Occupancy of energy levels significantly above what would correspond to thermal energies produces a Fermi pressure from the electrons. This stabilizes the stellar structure for long times, as it is independent of thermodynamic evolution and merely a result of the compactness. Radiative energy loss from the surface is modest due to their small surface, and provided by the heat capacity of the stellar matter as dominated by nucleons. The nucleons, on the other hand, thus are left at rather cold temperatures, and nuclear interactions do not contribute to energy generation. The dense nuclear matter of neutron stars is more complex, and the corresponding equation of state reflects the multi-nucleon interactions as they increase with compactness, and phase transitions to quark matter. Nuclear matter within a neutron star can be considered as ‘cold’ [38], and quark-gluon plasma aspects only arise at extreme compactness (unlike when neutron stars are formed during gravitational collapse of a star; see below). Radiative losses are modest due to the small surface of neutron stars, with typical radii of 10-15 km at masses of 1-2 M [364].
@@
Stellar structure and stability can be calculated from the relativistic description of behavior of such matter and using its equation of state, the Tolman-Oppenheimer-Volkoff equations [362, 469]. Neutron stars turn out to be excellent laboratories to study dense nuclear matter at two extremes of parameters that do not exist in nuclei [288, 458] (see also our discussion of Extremes of Matter in section 3.3). The neutron richness, or proton poorness, is extreme in neutron stars, the characteristic $Y_{e}$ variable of net charge and measuring deviations from neutron-proton balance ($Y_{e}=0.5$). The symmetry term of nuclear binding, hence, adopts extremes in neutron stars, and its connection to studies of nuclei with different neutron excess is a valuable diagnostic. The equation of state of neutron star matter describes its pressure dependency on thermodynamic variables of temperature and density. It shapes the compressibility of neutron star matter, which, on the other hand, determines tidal deformability of neutron stars as they collide in neutron star mergers, as well as the rebounce from gravitational collapse of a massive star as it forms a central neutron star, and develops an explosion or not [see review by 374, in Supernova Handbook]. The determination of the equation of state for nuclear matter throughout the range of density and temperature is one of the major challenges of nuclear astrophysics, involving nuclear theory as well as experiments using high-energy collisions of heavy ions, and astronomical observations of neutron star properties [38, 288].

### v2

@@ 1 Introduction
The field of Nuclear Astrophysics is situated between the two disciplines of nuclear physics and of astrophysics. Each of these includes experimental/observational and theoretical branches. The foundations of this field have been laid out in pioneering papers by Burbidge et al [57] and Cameron [60, 59], and later written up in textbooks such as Donald Clayton’s “Principles of stellar evolution and nucleosynthesis” [80], then “Cauldrons in the Cosmos” by Claus Rolfs and William Rodney [410], and the more-recent “Nuclear physics of stars” by Christian Iliadis [217] and “The synthesis of the elements” by Giora Shaviv [435]. Each of these take a different look at the complex interplay that occurs between the microscopic quantum processes that govern the nuclear reaction processes and the macroscopic processes of transporting energy and materials that characterize the conditions in stellar interiors. This results in considerable uncertainty for the analysis and determination of nuclear reaction processes in cosmic sites. On the other hand we understand enough about the origin of chemical elements to describe and teach this astrophysical and nuclear-physics problem [see above-cited textbooks and, for example, 57, 81, 495, 202, 24, 258, 89, 21, and other review articles that have been written over seven decades]). Often herein the descriptions of cosmic nucleosynthesis processes are inter-woven with our plausible understanding of its cosmic sources, thus mixing better consolidated with more-uncertain aspects. For astrophysicists, a description of cosmic sources which includes nuclear transformations remains a challenge, while nuclear physicists are challenged to determine nuclear reaction processes at conditions that are typically not available in the laboratory. The data base for this science are the measured outcomes of nuclear-reaction experiments, and the observations of isotopic variety and abundances in material samples or objects throughout the universe. Combining the characteristics of nuclear reactions with those of cosmic sites and their dynamics, nuclear astrophysics exploits astronomical data in attempts to rule out the invalid components of the complex and nested modeling that represents our understanding of the compositional evolution of matter in the universe.
@@
Conferences and workshops are of great importance in such a diverse field of nuclear astrophysics. They are exemplified by the series of conferences of "Nuclei in the Cosmos" (since 1990, no. XVIII held recently111see https://indico.icc.ub.edu/event/341/ for the most-recent NIC in 2025) and of "Nuclear Physics in Astrophysics" (since 2003, no. XI held recently222see https://events.hifis.net/event/540/ for the most recent NPA in 2024), typically held every 2 years. Gathering 200 people or more, these conferences may appear to novices as either overloaded with specialist discussions in sub-fields, or too shallow to address physics issues of interest with the required depth. Yet, they are essential to mediate among sub-fields, and are complemented by workshops addressing specific topics and questions in the field. International community networking projects such as IReNA (USA and international), JINA/CENAM (USA), ChETEC/ChETEC-Infra (EU/UK), UKAKUREN/JaFNA (Japan) and INAC (China) provide an umbrella for the diverse projects and for exchanges of tools; these have an important role to maintain the exchanges among diverse research groups. The experimental challenges ahead for this field are addressed by white papers that lobby for this science with the various funding agencies, see for example Lewitowicz and Widmann [293] and Schatz et al [428]. Nuclear astrophysics as a whole appears too broad a field to be addressed in a single review paper, due to this interdisciplinary nature. Partial reviews [e.g. 521] are useful to help appreciating the interplay between achievements, uncertainties, and tacit assumptions in each of the aspects of nuclear aspects.
@@ 2.1 General considerations
Only a subset of reactions, anticipated for a specific environment, are typically considered in astrophysical models. For hydrogen burning, the different branches of the pp chains [418] and the cyclic pattern of the four CNO cycles [518] are an early example of identifying reaction sequences that plausibly operate in stars. Similarly, the nuclear-reaction paths that may occur in cosmic sites are commonly organized in processes, or reaction patterns describing a reaction flow. These share a common environmental characteristic such as temperatures, densities, and abundance of reaction agents, and link a specific set of reactions. Regimes of reactions that are expected to occur under such characteristic conditions are illustrated in Figure 1 by white arrows, the processes of nucleosynthesis that are most-commonly discussed. These have been presented in the early period of nuclear astrophysics [57], and later a variety of other such processes and paths were proposed from different reasonings and scenarios [e.g. 197, 494, 531, 213, 529, 210, 426, 395, 379, 500, 381, 74, 75, (these will be discussed below in their respective context)]. Shown are the sequences of $\alpha$ capture reactions ($\alpha$ process) that may illustrate production of 56Ni and iron group elements from light nuclei, the continuation of heavier-nuclei production through slow neutron capture / $\beta$ decay sequences (s process) along the valley of stable nuclei (black squares), the alternate neutron capture path of rapid neutron captures (r process) far on the neutron-rich side, a rapid proton capture path (rp process) on the proton-rich side, and a horizontal production of proton-rich nuclei from stability through a p or $\gamma$ process, i.e. triggered by proton releases from ($\alpha$,p) or ($\gamma$,p) reactions on neutron-deficient oxygen or neon isotopes. Not shown are the generic $\beta$-decay reaction flows that drive unstable isotopes towards the valley of stability (i.e., forming diagonal lines in this Figure), driven only by the $\beta$-decay lifetimes of each isotope and thus (largely) independent from environmental characteristics. More such processes are being discussed in nuclear astrophysics as one aims to characterize a coherent nuclear-reaction network that may operate in specific cosmic sources. So, an intermittent n process [47, 381] and the intermediate i process [61, 74] are discussed to describe reactions at neutron intensities between the extremes characterizing the r and s process. Then, a $\nu$ process [531, 440] with its variants is used to characterize neutrino-driven nuclear reactions at high neutrino fluxes as they occur in the cores of gravitational collapse supernovae, and may characteristically alter the nucleosynthesis to produce specific rare isotopes that cannot be understood from the neutron capture nor $\gamma$ processes, such as 180Ta, 138La, or 19F. Also, reaction flows with electron captures and successive $\beta$ decays or vice versa through URCA pairs of nuclei define an URCA process [35]; not altering the composition, the URCA process establishes an energy leak through neutrino production, which affects burning-site conditions. These latter processes illustrate that details of nuclear reactions in cosmic sites are often more complex, and cannot all be shown in this Figure for clarity. This overview Figure 1 provides a valuable tool, to first order, for such explorations in order to find the most-relevant regions where nuclear-reaction properties matter for cosmic nucleosynthesis. It therefore is a key figure for the communication between nuclear physicists and astrophysicists.
@@
Uncertainties on reduced Coulomb repulsion in nuclear reactions under cosmic environments result from screening of the nuclear charge because of free electrons clustering around the positive charges. Particle beams in accelerator experiments are fundamentally different in their electron environments of the reaction compared to mixtures of high-temperature plasma in cosmic sites. Screening corrections are required for each of these two cases, and are different [see review by 13]. In the laboratory, potential extremes related to the screening towards the lowest energies motivate experiments studying reactions below keV energies between light nuclei such as 3He and deuterium towards 4He fusion [66, 489, 522].
@@
Experimental studies of nuclear physics for astrophysics include many indirect techniques, such as breaking up nuclei in heavy-ion collisions, or measuring surrogate reactions with similarities to the reactions and nuclear characteristics in question [see 521, for details].
@@ 2.2.1 Experimental Techniques and Methods
After the build-up, the subsequent 12C($\alpha,\gamma$)16O reaction in stellar helium burning is often called the holy grail reaction because it determines the cosmic abundance ratio between the two isotopes that are key for organic life, 12C and 16O. Numerous measurements have been performed to determine the reaction cross section towards lower energies. However, the interpretation of the data for a reliable extrapolation towards the Gamow range around 300 keV remains a challenge. The main difficulty is in the treatment of the pronounced interference effects between the low-energy tails of high-energy resonances, direct capture contributions and the high-energy tails of sub-threshold states. More specifically, the interpretation of the reaction cross-section [see 105, 104, for recent reviews] is based on contributions from a number of broad interfering $1^{-}$ and $2^{+}$ resonances in the compound nucleus $^{16}O$ as well as an E2 direct capture to the ground state of $^{16}O$. The $\gamma$ transition therefore shows an a E1 or a E2 multipolarity, with both components adding to the total cross section. Both components are predicted to have comparable strength due to the suppression of the E1 term by the breaking of isospin symmetry by Coulomb interaction as discussed by deBoer et al [104] To determine the cross-section data and map the resonance contributions interfering with bound state components extensive angular distribution in 12C($\alpha,\gamma$)16O experiments have been performed over a wide energy range.
@@ 2.5 Beta decay and neutrino reactions
In this context, the process of neutrino oscillations may alter the neutrino flavors between the neutrino source and the region where interactions with nuclei occur. This has been recognized and studied in recent years [e.g. 256] [see recent review by 146].
@@ 2.6 Nuclear theory complements
The original description of nuclear binding energies (through determining their masses) was made as an ensemble (macroscopic) effect through the liquid-drop model [515]. Since then, there had been an impressive evolution of theories [see 344, for a personal review]. A bottom-up approach starting from the nucleon-nucleon forces (microscopic approach) was pursued, in parallel to efforts aiming at an establishment of an empiric energy density functional (macroscopic approach) [504, 415]. Nuclear mass models are a compromise between empirical descriptions of macroscopic (and more-easily measurable) components and microscopic descriptions of nucleon-nucleon interactions deriving from fundamental theory or from empirical prescriptions. Nuclear theory attempts to advance models in specific regions of interest, while astrophysical nucleosynthesis models require consistent treatment of neutron captures and $\beta$ decays across the entire path of reactions; therefore, the FRDM baseline from 1995 is still widely used, as $\beta$ decays have been evaluated for a large number of nuclei here. Nuclear mass models evolved from the macroscopic finite range liquid drop model (FRDM) [347] to Hartree-Fock Bogoliubov (HFB) model versions of different sophistication levels, adding corrections to nucleon-nucleon forces with shell and deformation effects to a Fermi gas model for nuclei, among others [see 184, for testing model alternatives]. Towards the higher range of nuclear masses, nuclear deformations play a large role. The additional degrees of freedom for nuclear rotation and oscillations provide a larger phase space, thus making predictions of effective nuclear binding quite uncertain. Experimental results are analyzed and fitted to descriptions of mass models, thus determining model parameters; often models have a large number ($\simeq$ 30) of parameters determined in this way. As a measure, the 2-neutron separation energies $S_{2n}$ between models and measurements (where available) still show an RMS difference of 800 keV (see Figure 5), while reliable astrophysical r-process calculations would require a precision of order 100 keV [211]. Weak-interaction rates can be estimated theoretically in the QRPA approach [422]. But correspondingly, also the weak-interaction processes in atomic nuclei ($\beta$-decay rates, radioactive lifetimes) cannot be reliably calculated in theoretical models, if the details of shells and nuclear structure are not covered; shell models have proven to provide more realistic rates, in better agreement with experimental values [see 281, 187, for discussions of weak interactions in cosmic environments]. As agreements with experimental values are unsatisfactory, phenomenological models are widely used. This also extends to the determination of fission barriers, correspondingly [134, 345].
@@ 3.1 Stellar interiors
Once hydrogen in the core becomes exhausted, gravitational contraction results in a temperature increase until the next nuclear fuel 4He can undergo fusion reactions to liberate nuclear binding energy again. At the same time, hydrogen burning from then on proceeds in a shell [^5] Note that only $\simeq$ 10% of the total hydrogen will have been consumed during the main-sequence evolution., continuing throughout evolution in parallel to the nuclear processes in the stellar core (see Figure 7). The radiation pressure caused by nuclear energy release in the shell causes expansion of the outer zones of the star, while the core contracts due to a lack of energy production, removing the required thermal pressure for hydrostatic stability, and thus further expanding the envelope as a reaction to the contraction of the core. This is called the giant phase in stellar evolution [see 80, for a textbook introduction to stellar evolution]. Due to the smaller energy gain per 12C produced (7.3 MeV), the total helium-burning time is not quite as long as the main-sequence time; for the case of our Sun it is $\simeq$ 10% of the main sequence phase. This seems sufficiently long for most stars to still achieve and maintain hydrostatic equilibrium, rearranging temperatures and densities to the convective and radiative energy transport processes within the star and the nuclear energy gain in the core balancing radiative and wind losses at the surface.
@@
The presupernova evolution of massive stars in their late phases is a challenging topic in astrophysics. The above-discussed processes of mixing determine how evolution proceeds and modifies the structure of the star and its hydrostatic balance [see 517, for a recent broad-impact study]. Many different scenarios can be inferred from simulations, as mixing prescriptions and other parameters are varied. Examples are: burning shells progressing further inwards (instead of the more-plausible trend to proceed along the availability of unburnt fuel at larger distance from the core) further inwards with time [236, 238], or merge [407, 222], or hot-bottom burning [250]. Even though agreement is reached on the fundamental microscopic processes, their implementation in different numerical codes may provide different results, depending on how these processes are approximated and numerically stabilized within each code. Figure 7 illustrates an example; alternative codes for stellar evolution are maintained, e.g., by the Geneva [424, 130, 174, 338, 191] and Rome [71, 72, 296] groups. Astronomical observations may provide clarifications of such issues concerning stellar interiors (see Section 4 on Nuclear Astronomy below).
@@
Stars are often formed in groups, and multiple stars may be created as single entities yet close enough to have significant mutual interactions that affects their evolution. The impact of a companion star on stellar evolution had been described long ago [216]. The rather obvious effect of a companion to enhance mass loss and thus stripping the envelope has been modeled [533, 388, 467]. This is understood as a path towards explaining core-collapse supernovae which do not show a hydrogen envelope, and are thus observationally known as supernova types Ib and Ic. More difficult are successive periods of mass exchange in different directions within a binary system, or phases of a common envelope [e.g. 508]. Thus, binary interactions and their impacts have developed into a major field of stellar evolution [216, 484, 388, 421, 100, 225, 486, 283, 321, 429]. This is mainly addressed by theory and simulations, because it is difficult to obtain and exploit observational constraints on models of binary interactions, as observational phenomena cannot easily by attributed to binary interaction physics as discriminated from instabilities in single-star evolution.
@@
Compact stars are stabilized by degeneracy pressure from electrons (the white dwarf stars), and by the complex pressure behavior of dense nuclear matter called its equation of state (the neutron stars). White dwarfs are the compact remnants of stars after the giant stage and strong mass losses therein [see 423, for a recent review]. Typical mass and size values are 0.16 M and 0.01 R, respectively. Matter is confined to a rather small volume in compact stars. Electrons move freely as a Fermi gas within the charge-neutral star which is bound by gravity that is caused mainly by the nucleons. As thermal energies become small and approach zero for the electrons, their kinetic energy level occupancy obeys quantum constrains and the Pauli exclusion principle for identical states. Occupancy of energy levels significantly above what would correspond to thermal energies produces a Fermi pressure from the electrons. This stabilizes the stellar structure for long times, as it is independent of thermodynamic evolution and merely a result of the compactness. Radiative energy loss from the surface is modest due to their small surface, and provided by the heat capacity of the stellar matter as dominated by nucleons. The nucleons, on the other hand, thus are left at rather cold temperatures, and nuclear interactions do not contribute to energy generation. The dense nuclear matter of neutron stars is more complex, and the corresponding equation of state reflects the multi-nucleon interactions as they increase with compactness, and phase transitions to quark matter. Nuclear matter within a neutron star can be considered as ‘cold’ [39], and quark-gluon plasma aspects only arise at extreme compactness (unlike when neutron stars are formed during gravitational collapse of a star; see below). Radiative losses are modest due to the small surface of neutron stars, with typical radii of 10-15 km at masses of 1-2 M [367].
@@
Stellar structure and stability can be calculated from the relativistic description of behavior of such matter and using its equation of state, the Tolman-Oppenheimer-Volkoff equations [365, 473]. Neutron stars turn out to be excellent laboratories to study dense nuclear matter at two extremes of parameters that do not exist in nuclei [289, 461] (see also our discussion of Extremes of Matter in section 3.3). The neutron richness, or proton poorness, is extreme in neutron stars, the characteristic $Y_{e}$ variable of net charge and measuring deviations from neutron-proton balance ($Y_{e}=0.5$). The symmetry term of nuclear binding, hence, adopts extremes in neutron stars, and its connection to studies of nuclei with different neutron excess is a valuable diagnostic. The equation of state of neutron star matter describes its pressure dependency on thermodynamic variables of temperature and density. It shapes the compressibility of neutron star matter, which, on the other hand, determines tidal deformability of neutron stars as they collide in neutron star mergers, as well as the rebounce from gravitational collapse of a massive star as it forms a central neutron star, and develops an explosion or not [see review by 377, in Supernova Handbook]. The determination of the equation of state for nuclear matter throughout the range of density and temperature is one of the major challenges of nuclear astrophysics, involving nuclear theory as well as experiments using high-energy collisions of heavy ions, and astronomical observations of neutron star properties [39, 289].